# Full-sky Models of Galactic Microwave Emission and Polarization at Sub-arcminute Scales for the Python Sky Model

### v1

@@ Abstract
Polarized foreground emission from the Galaxy is one of the biggest challenges facing current and upcoming cosmic microwave background (CMB) polarization experiments. We develop new models of polarized Galactic dust and synchrotron emission at CMB frequencies that draw on the latest observational constraints, that employ the “polarization fraction tensor” framework to couple intensity and polarization in a physically motivated way, and that allow for stochastic realizations of small-scale structure at sub-arcminute angular scales currently unconstrained by full-sky data. We implement these models into the publicly available Python Sky Model (PySM) software and additionally provide PySM interfaces to select models of dust and CO emission from the literature. We characterize the behavior of each model by quantitatively comparing it to observational constraints in both maps and power spectra, demonstrating an overall improvement over previous PySM models. Finally, we synthesize models of the various Galactic foreground components into a coherent suite of three plausible microwave skies that span a range of astrophysical complexity allowed by current data. [^1] A supplement describing author contributions to this paper can be found at https://pysm3.readthedocs.io/en/latest/pysm_methods_author_contributions.html.
software: Astropy (Astropy Collaboration et al., 2013, 2018, 2022), cmocean (Thyng et al., 2016), HEALPix (Górski et al., 2005), healpy (Zonca et al., 2019), Matplotlib (Hunter, 2007), NaMaster (Alonso et al., 2019), NumPy (van der Walt et al., 2011; Harris et al., 2020), Pynkowski [^19] https://javicarron.github.io/pynkowski/pynkowski.html, PySM (Thorne et al., 2017; Zonca et al., 2021), SciPy (Virtanen et al., 2020)
1 Introduction
One of the principal challenges for current and future cosmic microwave background (CMB) polarization experiments is mitigating contaminating emission from the Galaxy. Polarized Galactic emission is brighter than current upper limits on a primordial $B$-mode signal at all frequencies, even in particularly clean patches of the high Galactic latitude sky (Planck Collaboration XI, 2020; Ade et al., 2021). Many ongoing and upcoming surveys such as the Simons Observatory (Ade et al., 2019), CMB-S4 (Abazajian et al., 2022), and LiteBIRD (LiteBIRD Collaboration et al., 2023) observe large sky fractions and will need to mitigate foreground emission potentially many times brighter than the targeted cosmological signals. Understanding the potential complexities of Galactic emission and designing analyses that are robust to these complexities are of paramount importance for constraining the physics of the early Universe.
@@ 1 Introduction
To this end, tools have been developed to simulate full-sky, multi-frequency realizations of Galactic emission drawing both on data-driven constraints and theoretical models. The Planck Sky Model (PSM; Delabrouille et al., 2013) was originally built to develop data analysis tools for the Planck mission, using pre-existing data sets. The PSM evolved throughout the Planck mission timeline to include Planck observations and adjust to Planck data analysis requirements, and it has been widely used in various data challenges and for planning future CMB and 21 cm line mapping experiments (e.g., Remazeilles et al., 2018; Fornazier et al., 2022; Ghosh et al., 2022). Building on the PSM, the Python Sky Model (PySM; Thorne et al., 2017) provides a Python interface to expanded classes of foreground models. PySM has been employed for both forecasting (e.g., Abazajian et al., 2022; Hensley et al., 2022; CCAT-Prime Collaboration et al., 2023; Wolz et al., 2024) and data analysis (e.g., BICEP2 Collaboration et al., 2018; Vacher et al., 2023; Ade et al., 2025).
@@
We organize the paper as follows: Section 2 is an overview of how Galactic emission models are implemented in PySM; Section 3 presents our methodology for generating stochastic simulations of small-scale emission using the polarization fraction tensor formalism; Section 4 describes our implementation of an alternative dust model and of a suite of CO emission models; Section 5 details a collection of validation metrics for the new models; Section 6 discusses limitations of the models presented here and future directions for development; Section 7 presents our proposed suite of three sky models that span a range of complexity; and Section 8 concludes with a summary.
2 Modeling Framework
2.1 The PySM Software
@@
The PySM software [^2] https://pysm3.readthedocs.io/en/latest/index.html provides a convenient interface for generating full-sky maps of total and polarized emission at far-infrared through radio frequencies. Users can select one or more emission mechanisms to simulate, including the CMB, dust, synchrotron radiation, free-free emission, and CO line emission. Most emission mechanisms have more than one model to select from, where here a model refers both to the spatial morphology and frequency dependence of the emission. Stokes $I$, $Q$, and $U$ maps are provided in HEALPix [^3] https://healpix.sourceforge.io (Górski et al., 2005) format at the requested $N_{\rm side}$ at a user-specified frequency or integrated over a user-specified bandpass.
@@
One of the aims of this work is to extend the highest angular resolution of the maps that PySM can generate. The first and most stringent challenge of such high resolutions is the sheer size of the templates: a single $I$, $Q$, or $U$ map at a pixel size of $0.4\arcmin$ ($N_{\rm side}=8192$) occupies  3 GB in single precision and is 256 times larger than an original PySM 2 map. The groundwork that allowed the implementation of this new generation of Galactic models started in 2019 with the rewrite of PySM and its release as PySM 3 (see Zonca et al., 2021, for details). The problem of distribution has been solved by hosting all the input templates at NERSC [^4] https://portal.nersc.gov/project/cmb/pysm-data, with templates downloaded and cached by PySM as needed using the facilities included in astropy.data (Astropy Collaboration et al., 2013, 2018). Moreover, all the members of the CMB community using NERSC for computing are able to directly access the same folders locally.
@@
The second issue is memory usage. PySM 3 leverages numba (Lam et al., 2015) to compile Python on-the-fly to machine code so that the evaluation and bandpass integration of each model avoids the temporary arrays allocated by numpy and reduces memory consumption at least by a factor of two. Moreover, numba supports multi-threading so it can make use of all the cores available in the system. Thanks to these improvements, foreground models with a resolution up to $N_{\rm side}=8192$ can be generated while keeping the disk, memory, and CPU requirements manageable. We defer to Section 3 a discussion of the algorithmic improvements that enable models to be constructed at these small angular scales.
2.2 Overview of Galactic Emission Mechanisms
@@
The Galactic interstellar medium (ISM) consists of matter and radiation in various forms. At the microwave wavelengths relevant for CMB observations there are several relevant emission mechanisms, each with their own frequency dependence, and different mechanisms dominate the emission from different regions of the ISM. The spatial distribution of microwave emission thus varies with frequency (e.g., Planck Collaboration X, 2016).
@@
Cold ($\sim$10–30 K) grains of interstellar dust emit thermal vibrational radiation with a spectrum that peaks at $\sim$2 THz. The dust emission spectrum shows an excess over expectations from thermal vibrational emission below $\nu=100$ GHz: this excess is dubbed anomalous microwave emission (AME). AME is thought to be electric dipole radiation from rapidly spinning, sub-nanometer dust grains (Draine & Lazarian, 1998a). Relativistic cosmic ray electrons spiraling in the Galactic magnetic field emit synchrotron radiation, while warm ionized gas emits free-free radiation (or bremsstrahlung) through the interaction of free electrons with ions. Synchrotron and free-free emission typically dominate the Galactic ISM emission at frequencies $\sim$10–100 GHz. Finally, atoms and molecules in the Galactic ISM, through vibrational and rotational shifts in energy levels, emit radiation in the form of a rich spectrum of discrete lines. Most notable among these is a bright comb of carbon monoxide (CO) emission lines at integer multiples of the $J=1\rightarrow 0$ ($\nu=115$ GHz) rotational transition.
@@
All of these emission mechanisms are capable of producing polarized radiation. The aggreg ate emission from aspherical dust grains aligned with their short axes preferentially parallel to the local magnetic field is polarized perpendicular to the projected field orientation at levels up to $\sim$20(Planck Collaboration Int. XIX, 2015). Synchrotron emission is inherently polarized at the $\sim$75level and, like dust emission, is polarized perpendicular to the projection of the local magnetic field onto the plane of the sky (Rybicki & Lightman, 1986). Dust and synchrotron polarization have both been robustly detected and are the principal polarized Galactic foregrounds at microwave frequencies. The relevance of the other mechanisms is less clear. CO line emission can be linearly polarized via the Goldreich-Kylafis effect (Goldreich & Kylafis, 1981), though there have been few detections to date (e.g., Greaves et al., 1999, 2002; Cortes et al., 2008; Houde et al., 2013). If the sub-nanometer dust grains believed to be responsible for the AME are able to align with the local magnetic field, then AME will be polarized (Draine & Lazarian, 1998b). Searches for AME polarization in specific objects and over large sky areas have so far yielded only upper limits (e.g., Génova-Santos et al., 2017; Herman et al., 2023). Finally, free-free emission is inherently unpolarized, but a small level of polarization can be produced near the edges of H II regions from Thomson scattering (Rybicki & Lightman, 1986). However, this effect is relevant only at much higher angular resolution than typically employed in CMB observations.
@@ 2.2.1 Dust Emission
where $S$ is any of $I$, $Q$, or $U$ and $B_{\nu}\left(T\right)$ is the Planck function. The parameter $A_{d}^{S}$ is the dust intensity $S_{\nu}$ at the reference frequency $\nu_{0}$, here taken to be 353 GHz. We refer to the $A_{d}^{S}$ as “amplitude” parameters. The parameters $\beta_{d}$ and dust temperature $T_{d}$ describe the frequency dependence of the emission. We refer to them as “spectral” parameters. Typical values of the spectral parameters for dust emission in both total intensity and polarization are $\beta_{d}=1.5$ and $T_{d}=20$ K (Planck Collaboration XI, 2020), with variations of order 10% observed in total intensity in the high Galactic latitude sky (e.g., Planck Collaboration X, 2016; Planck Collaboration Int. XLVIII, 2016).
In this work, we present the new dust models d9, d10, d11 (Section 3) and an implementation of the Martínez-Solaeche et al. (2018) (“MKD”) model as d12 (Section 4.1). We employ previous PySM dust models only for purposes of comparison.
2.2.2 Synchrotron Emission
@@
where $S$ is any of $I$, $Q$, or $U$. The parameter $A_{s}^{S}$ is the synchrotron intensity $S_{\nu}$ at the reference frequency $\nu_{0}$, here taken to be 23 GHz. As with dust, we refer to the $A_{s}$ as “amplitude” parameters. The parameters $\beta_{s}$, $c_{s}$, and $\nu_{c}$ describe the frequency dependence of the emission. In all models, we take $\nu_{c}=23$ GHz. When expressing $S_{\nu}$ in specific intensity units, $\beta_{s}\simeq-1$.
@@
In this work, we present the new synchrotron models s5, s6, and s7 (Section 3). We employ previous PySM synchrotron models only for purposes of comparison.
2.2.3 Free-free Emission
@@
We do not develop new models of free-free emission in this work, but rather rely on the existing f1 model (Thorne et al., 2017). The frequency dependence of free-free emission is known from theory (Draine, 2011, and references therein), but over the range of frequencies relevant for microwave observations, it can be approximated as a simple power law. The f1 model assumes a sky-constant power-law behavior $S_{\nu}\propto\nu^{-0.14}$ in specific intensity units. The amplitude is given by the free-free amplitude map from the Commander component separation analysis (Planck Collaboration X, 2016), which has an angular resolution of $1^{\circ}$. At angular scales $<1^{\circ}$, the emission is based on Gaussian random fluctuations modulated by the intensity map (see Thorne et al., 2017, for details). Free-free emission is assumed to be unpolarized.
2.2.4 Anomalous Microwave Emission (AME)
@@
We do not develop new models of AME in this work, but rather rely on the existing a1 and a2 models (Thorne et al., 2017). These models are based on a Commander component separation analysis that produced full-sky maps of AME amplitude and spectral parameters at $1^{\circ}$ resolution (Planck Collaboration X, 2016). The AME frequency dependence is modeled as a sum of two components each having spectra based on theoretical models computed by the SpDust software (Ali-Haïmoud et al., 2009; Silsbee et al., 2011). While each component is described by an amplitude and a peak frequency, one component was required to have a sky-constant peak frequency, fit to a value of 33.35 GHz (Planck Collaboration X, 2016). Thus, the AME model is based upon three maps: the amplitude of each of the two components and the peak frequency of one component. Emission at scales $<1^{\circ}$ is generated using the higher-resolution observations of thermal dust emission, where the spatially varying scaling factor is determined by convolving the thermal dust emission map to $1^{\circ}$ resolution and taking the ratio with the AME map (Thorne et al., 2017).
@@
The a1 and a2 models differ only in their treatment of polarization. AME polarization has not been detected, and recent analysis places an upper limit on the intrinsic polarization fraction of $\lesssim 3$% (Herman et al., 2023). The a1 model assumes that AME is unpolarized, whereas the a2 model assumes a constant polarization fraction of 2%. The polarization angle of the AME in the a2 model is based on the polarization angle of the 353 GHz polarized dust emission determined by component separation with Commander (Planck Collaboration X, 2016). As with total intensity, the a2 AME polarization map inherits the small-scale fluctuations introduced to the thermal dust polarization map (see Thorne et al., 2017, for details).
2.2.5 CO Emission

### v2

@@ Abstract
Polarized foreground emission from the Galaxy is one of the biggest challenges facing current and upcoming cosmic microwave background (CMB) polarization experiments. We develop new models of polarized Galactic dust and synchrotron emission at CMB frequencies that draw on the latest observational constraints, that employ the “polarization fraction tensor” framework to couple intensity and polarization in a physically motivated way, and that allow for stochastic realizations of small-scale structure at sub-arcminute angular scales currently unconstrained by full-sky data. We implement these models into the publicly available Python Sky Model (PySM) software and additionally provide PySM interfaces to select models of dust and CO emission from the literature. We characterize the behavior of each model by quantitatively comparing it to observational constraints in both maps and power spectra, demonstrating an overall improvement over previous PySM models. Finally, we synthesize models of the various Galactic foreground components into a coherent suite of three plausible microwave skies that span a range of astrophysical complexity allowed by current data. [^1] A supplement describing author contributions to this paper can
be found at https://pysm3.readthedocs.io/en/latest/pysm_methods_author_contributions.html.
software: Astropy (AstropyCollaboration:2013; AstropyCollaboration:2018; AstropyCollaboration:2022), cmocean (Thyng:2016), HEALPix (Gorski:2005), healpy (Zonca:2019), Matplotlib (Hunter:2007), NaMaster (Alonso:2019), NumPy (vanderWalt:2011; Harris:2020), Pynkowski [^19] https://javicarron.github.io/pynkowski/pynkowski.html, PySM (Thorne:2017; Zonca:2021), SciPy (Virtanen:2020)
I Introduction
One of the principal challenges for current and future cosmic microwave background (CMB) polarization experiments is mitigating contaminating emission from the Galaxy. Polarized Galactic emission is brighter than current upper limits on a primordial $B$-mode signal at all frequencies, even in particularly clean patches of the high Galactic latitude sky (planck2016-l11A; Ade:2021). Many ongoing and upcoming surveys such as the Simons Observatory (Ade:2019), CMB-S4 (Abazajian:2022), and LiteBIRD (LiteBIRDCollaboration:2023) observe large sky fractions and will need to mitigate foreground emission potentially many times brighter than the targeted cosmological signals. Understanding the potential complexities of Galactic emission and designing analyses that are robust to these complexities are of paramount importance for constraining the physics of the early Universe.
@@ I Introduction
To this end, tools have been developed to simulate full-sky, multi-frequency realizations of Galactic emission drawing both on data-driven constraints and theoretical models. The Planck Sky Model (PSM; delabrouille2012) was originally built to develop data analysis tools for the Planck mission, using pre-existing data sets. The PSM evolved throughout the Planck mission timeline to include Planck observations and adjust to Planck data analysis requirements, and it has been widely used in various data challenges and for planning future CMB and 21 cm line mapping experiments (e.g., Remazeilles:2018; Fornazier:2022; Ghosh:2022). Building on the PSM, the Python Sky Model (PySM; Thorne:2017) provides a Python interface to expanded classes of foreground models. PySM has been employed for both forecasting (e.g., Abazajian:2022; Hensley:2022; CCAT-PrimeCollaboration:2023; Wolz:2024) and data analysis (e.g., BICEP2Collaboration:2018; Vacher:2023; Ade:2025b).
@@
We organize the paper as follows: Section II is an overview of how Galactic emission models are implemented in PySM; Section III presents our methodology for generating stochastic simulations of small-scale emission using the polarization fraction tensor formalism; Section IV describes our implementation of an alternative dust model and of a suite of CO emission models; Section V details a collection of validation metrics for the new models; Section VI discusses limitations of the models presented here and future directions for development; Section VII presents our proposed suite of three sky models that span a range of complexity; and Section VIII concludes with a summary.
II Modeling Framework
II.1 The PySM Software
@@
The PySM software [^2] https://pysm3.readthedocs.io/en/latest/index.html provides a convenient interface for generating full-sky maps of total and polarized emission at far-infrared through radio frequencies. Users can select one or more emission mechanisms to simulate, including the CMB, dust, synchrotron radiation, free-free emission, and CO line emission. Most emission mechanisms have more than one model to select from, where here a model refers both to the spatial morphology and frequency dependence of the emission. Stokes $I$, $Q$, and $U$ maps are provided in HEALPix [^3] https://healpix.sourceforge.io (Gorski:2005) format at the requested $N_{\rm side}$ at a user-specified frequency or integrated over a user-specified bandpass.
@@
One of the aims of this work is to extend the highest angular resolution of the maps that PySM can generate. The first and most stringent challenge of such high resolutions is the sheer size of the templates: a single $I$, $Q$, or $U$ map at a pixel size of $0.4\arcmin$ ($N_{\rm side}=8192$) occupies  3 GB in single precision and is 256 times larger than an original PySM 2 map. The groundwork that allowed the implementation of this new generation of Galactic models started in 2019 with the rewrite of PySM and its release as PySM 3 (see Zonca:2021, for details). The problem of distribution has been solved by hosting all the input templates at NERSC [^4] https://portal.nersc.gov/project/cmb/pysm-data, with templates downloaded and cached by PySM as needed using the facilities included in astropy.data (AstropyCollaboration:2013; AstropyCollaboration:2018). Moreover, all the members of the CMB community using NERSC for computing are able to directly access the same folders locally.
@@
The second issue is memory usage. PySM 3 leverages numba (Lam:2015) to compile Python on-the-fly to machine code so that the evaluation and bandpass integration of each model avoids the temporary arrays allocated by numpy and reduces memory consumption at least by a factor of two. Moreover, numba supports multi-threading so it can make use of all the cores available in the system. Thanks to these improvements, foreground models with a resolution up to $N_{\rm side}=8192$ can be generated while keeping the disk, memory, and CPU requirements manageable. We defer to Section III a discussion of the algorithmic improvements that enable models to be constructed at these small angular scales.
II.2 Overview of Galactic Emission Mechanisms
@@
The Galactic interstellar medium (ISM) consists of matter and radiation in various forms. At the microwave wavelengths relevant for CMB observations there are several relevant emission mechanisms, each with their own frequency dependence, and different mechanisms dominate the emission from different regions of the ISM. The spatial distribution of microwave emission thus varies with frequency (e.g., planck2014-a12).
@@
Cold ($\sim$10–30 K) grains of interstellar dust emit thermal vibrational radiation with a spectrum that peaks at $\sim$2 THz. The dust emission spectrum shows an excess over expectations from thermal vibrational emission below $\nu=100$ GHz: this excess is dubbed anomalous microwave emission (AME). AME is thought to be electric dipole radiation from rapidly spinning, sub-nanometer dust grains (Draine:1998b). Relativistic cosmic ray electrons spiraling in the Galactic magnetic field emit synchrotron radiation, while warm ionized gas emits free-free radiation (or bremsstrahlung) through the interaction of free electrons with ions. Synchrotron and free-free emission typically dominate the Galactic ISM emission at frequencies $\sim$10–100 GHz. Finally, atoms and molecules in the Galactic ISM, through vibrational and rotational shifts in energy levels, emit radiation in the form of a rich spectrum of discrete lines. Most notable among these is a bright comb of carbon monoxide (CO) emission lines at integer multiples of the $J=1\rightarrow 0$ ($\nu=115$ GHz) rotational transition.
All of these emission mechanisms are capable of producing polarized radiation. Dust and synchrotron polarization have both been robustly detected and are the principal polarized Galactic foregrounds at microwave frequencies. Dust grains rotate about their short axis—the axis of greatest moment of inertia—and the rotation axis in turn aligns with the local magnetic field due to the magnetic moment of the grain. The aggregate emission from a population of such grains is polarized perpendicular to the projected field orientation at levels up to $\sim$20% (planck2014-XIX). Synchrotron emission is inherently polarized at the $\sim$75% level and, like dust emission, is polarized perpendicular to the projection of the local magnetic field onto the plane of the sky (Rybicki:1986).
@@
The relevance of the other mechanisms is less clear. CO line emission can be linearly polarized via the Goldreich-Kylafis effect (Goldreich:1981), though there have been few detections to date (e.g., Greaves:1999; Greaves:2002; Cortes:2008; Houde:2013). If the sub-nanometer dust grains believed to be responsible for the AME are able to align with the local magnetic field, then AME will be polarized (Draine:1998a). Searches for AME polarization in specific objects and over large sky areas have so far yielded only upper limits (e.g., Genova-Santos:2017; Herman:2023). Finally, free-free emission is inherently unpolarized, but a small level of polarization can be produced near the edges of H II regions from Thomson scattering (Rybicki:1986). However, this effect is relevant only at much higher angular resolution than typically employed in CMB observations.
@@ II.2.1 Dust Emission
where $S$ is any of $I$, $Q$, or $U$ and $B_{\nu}\left(T\right)$ is the Planck function. The parameter $A_{d}^{S}$ is the dust intensity $S_{\nu}$ at the reference frequency $\nu_{0}$, here taken to be 353 GHz. We refer to the $A_{d}^{S}$ as “amplitude” parameters, with each pixel in a map having its own $A_{d}$ value.
The parameters $\beta_{d}$ and dust temperature $T_{d}$ describe the frequency dependence of the emission. We refer to them as “spectral” parameters. Typical values of the spectral parameters for dust emission in both total intensity and polarization are $\beta_{d}=1.5$ and $T_{d}=20$ K (planck2016-l11A), with variations of order 10% observed in total intensity in the high Galactic latitude sky (e.g., planck2014-a12; planck2016-XLVIII).
In this work, we present the new dust models d9, d10, d11 (Section III) and an implementation of the Martinez-Solaeche:2018 (“MKD”) model as d12 (Section IV.1). We employ previous PySM dust models only for purposes of comparison.
II.2.2 Synchrotron Emission
@@
where $S$ is any of $I$, $Q$, or $U$. The parameter $A_{s}^{S}$ is the synchrotron intensity $S_{\nu}$ at the reference frequency $\nu_{0}$, here taken to be 23 GHz. As with dust, we refer to the $A_{s}$ as “amplitude” parameters that vary from pixel to pixel in a map.
The parameters $\beta_{s}$, $c_{s}$, and $\nu_{c}$ describe the frequency dependence of the emission. Nonzero $c_{s}$ has been found in both simulations of Galactic synchrotron emission (Jaffe:2011) and in observational data (Kogut:2012). It arises from a non-power law energy distribution of the cosmic ray electrons responsible for the emission, due for instance to radiative energy losses. In all models, we take $\nu_{c}=23$ GHz. When expressing $S_{\nu}$ in specific intensity units, $\beta_{s}\simeq-1$.
@@
In this work, we present the new synchrotron models s5, s6, and s7 (Section III). We employ previous PySM synchrotron models only for purposes of comparison.
II.2.3 Free-free Emission
@@
We do not develop new models of free-free emission in this work, but rather rely on the existing f1 model (Thorne:2017). The frequency dependence of free-free emission is known from theory (Draine:2011, and references therein), but over the range of frequencies relevant for microwave observations, it can be approximated as a simple power law. The f1 model assumes a sky-constant power-law behavior $S_{\nu}\propto\nu^{-0.14}$ in specific intensity units. The amplitude is given by the free-free amplitude map from the Commander component separation analysis (planck2014-a12), which has an angular resolution of $1^{\circ}$. At angular scales $1^{\circ}$, the emission is based on Gaussian random fluctuations modulated by the intensity map (see Thorne:2017, for details). Free-free emission is assumed to be unpolarized.
II.2.4 Anomalous Microwave Emission (AME)
@@
We do not develop new models of AME in this work, but rather rely on the existing a1 and a2 models (Thorne:2017). These models are based on a Commander component separation analysis that produced full-sky maps of AME amplitude and spectral parameters at $1^{\circ}$ resolution (planck2014-a12). The AME frequency dependence is modeled as a sum of two components each having spectra based on theoretical models computed by the SpDust software (Ali-Haimoud:2009; Silsbee:2011). While each component is described by an amplitude and a peak frequency, one component was required to have a sky-constant peak frequency, fit to a value of 33.35 GHz (planck2014-a12). Thus, the AME model is based upon three maps: the amplitude of each of the two components and the peak frequency of one component. Emission at scales $1^{\circ}$ is generated using the higher-resolution observations of thermal dust emission, where the spatially varying scaling factor is determined by convolving the thermal dust emission map to $1^{\circ}$ resolution and taking the ratio with the AME map (Thorne:2017).
@@
The a1 and a2 models differ only in their treatment of polarization. AME polarization has not been detected, and recent analysis places an upper limit on the intrinsic polarization fraction of $\lesssim 3$% (Herman:2023). The a1 model assumes that AME is unpolarized, whereas the a2 model assumes a constant polarization fraction of 2%. The polarization angle of the AME in the a2 model is based on the polarization angle of the 353 GHz polarized dust emission determined by component separation with Commander (planck2014-a12). As with total intensity, the a2 AME polarization map inherits the small-scale fluctuations introduced to the thermal dust polarization map (see Thorne:2017, for details).
II.2.5 CO Emission